The wad of clay A is projected as shown at the same instant that cylinder B is released. The two bodies collide and stick together at C and then ultimately strike the horizontal surface at D. Determine the horizontal distance d. Use the values vo = 12 m/s, ? = 40°, L = 6 m, mA-0.1 kg, and mp 0.2 kg. - ?? Uo Problem 3/212
Answer #1

classmAte Date Vo 一之 t = 0:6527 sec. 0°)(0-6527) ー9.81 (0.6527)2 a= 2.945 m . Bc= h-a = 5,0346-2-345 -0896m 6.403 m/s mis

ClASSMAte Date Page In homizntrechon 6 = 3.064 (0.477) |d- 1-462- う m) nsuser